# The Mean-Field Dynamics of Transformers

### v1

@@ Abstract.
We develop a mathematical framework that interprets Transformer attention as an interacting particle system and studies its continuum (mean-field) limits. By idealizing attention continuous on the sphere, we connect Transformer dynamics to Wasserstein gradient flows, synchronization models (Kuramoto), and mean-shift clustering. Central to our results is a global clustering phenomenon whereby tokens cluster asymptotically after long metastable states where they are arranged into multiple clusters. We further analyze a tractable equiangular reduction to obtain exact clustering rates, show how commonly used normalization schemes alter contraction speeds, and identify a phase transition for long-context attention. The results highlight both the mechanisms that drive representation collapse and the regimes that preserve expressive, multi-cluster structure in deep attention architectures.
@@
A standard Transformer is a stack of layers, each consisting of two main components: a multi-head attention block and a feed-forward (MLP) block. Both components are followed by normalization and residual connections. Schematically, if $X\in\mathbb{R}^{n\times d}$ denotes the matrix of $n$ tokens of dimension $d$, a single layer performs the composition of the two following maps:
$X$ $\longmapsto X+\mathrm{Attention}(X),$
$X$ $\longmapsto X+\mathrm{MLP}(X).$
Normalization steps are inserted before or after each block depending on the implementation; see Section 6.2 for some examples. These residual connections, together with layer normalization [ba2016layer, zhang2019root], play a crucial stabilizing role, analogous to skip connections in residual networks [he2016identity, haber2017stable]. We will later see that such normalization can be idealized as a projection onto the unit sphere, a simplification that preserves essential dynamical features of trained models.
@@
$\mathrm{Attention}(X)=A_{\beta}(QX,KX)\,VX,\qquad A_{\beta}(QX,KX)_{ij}=\frac{\exp\!\big(\beta\langle(QX)_{i},(KX)_{j}\rangle\big)}{\sum_{k=1}^{n}\exp\!\big(\beta\langle(QX)_{i},(KX)_{k}\rangle\big)}.$ (1)
The stochastic matrix $A_{\beta}$ encodes the pairwise influence between tokens: its $(i,j)$-th entry represents the degree to which token $i$ attends to token $j$. Equation (1) can thus be interpreted as a nonlinear interaction rule in which each token updates as a weighted average of all others, with weights depending on their similarity in feature space. This operator will later serve as the starting point for the simplified continuous-time models studied in this paper.
@@
Modern variants of the Transformer introduce additional components: hierarchical structures for vision tasks [liu2021swin], sparse or long-range attention [beltagy2020longformer, child2019generating], low-rank adaptation mechanisms [hu2022lora], and mixture-of-experts [shazeer2017outrageously, boixadsera2025powerfinegrainedexpertsgranularity] architectures that route tokens through specialized subnetworks. These extensions have led to dramatic empirical gains but lie outside the scope of this paper. Our goal is to isolate and analyze the mathematical core of the architecture—the attention mechanism itself—and to understand how it regulates token dynamics.
@@
$\mathcal{E}_{\beta}(\mu)=\frac{1}{2}\iint e^{\beta\langle x,y\rangle}\,\mathrm{d}\mu(x)\mathrm{d}\mu(y).$ (3)
In the large-$\beta$ regime, this gradient-flow structure amplifies nonlinear diffusion effects and, at leading order, the continuity equation associated with (SA) converges formally to a reverse heat equation. This anti-diffusive limit foreshadows the clustering behavior described in the next section. In contrast, for (SA) with appropriate time rescaling, it converges to a porous medium type equation; see [bruno2025multiscale].
The (SA) and (USA) dynamics are minimal but faithful abstractions of self-attention. They retain its essential nonlinear and nonlocal features while remaining amenable to analysis. As discussed in the next section, both exhibit clustering effects similar to those observed in trained Transformers, providing a tractable framework to study attention-driven representation dynamics. In fact, when restricted to the circle, these dynamics reduce to variants of the Kuramoto model, the classical framework to study synchronization which correspond precisely to the clustering behavior alluded to above.
$3.$ The Kuramoto connection.
@@
We now turn to one of the most striking properties of the dynamics (USA) and (SA): the spontaneous emergence of clusters. In fact, clustering can be observed in trained Transformer models: Figure 1 indicates token embeddings exhibit a progressive concentration of pairwise inner products near 11, revealing the gradual formation of clusters.
Figure 1. Histograms of pairwise inner products $\{\langle x_{i}(t),x_{j}(t)\rangle\}_{i\neq j}$ at layers $t=0$, 55, 3030, and 4848 in the pre-trained ALBERT XLarge v2 model [lanalbert] available on Huggingface. The progressive concentration of mass near 11 across layers illustrates the emergence of clustering in token embeddings. Figure reproduced from [geshkovski2025mathematical].
@@
The result follows from the general theory of consensus dynamics on compact manifolds [markdahl2017almost]. Since the dynamics (SA) are a smooth gradient flow of an analytic energy functional, the classical theorem of Łojasiewicz [lojasiewicz1963propriete] guarantees that every trajectory converges to a stationary point. A detailed analysis of the critical set shows that, except for the fully synchronized configuration, all other stationary points are saddles by exhibiting an escape direction. This result also holds for weighted tokens that evolve according to the Wasserstein-Fisher-Rao gradient flow whereby particles evolve and are also dynamically reweighted [chen2025clustering]. By the center-stable manifold theorem [shub2013global, Thm. III.7], the set of initial conditions whose trajectories converge to such saddles is contained in a countable union of lower-dimensional manifolds and thus has measure zero. Consequently, almost every initialization leads to convergence toward the synchronized equilibrium. This argument, first established in [markdahl2017almost] and subsequently refined in [criscitiello2024synchronization], provides a complete characterization of the long-time behavior of the (USA) dynamics—see [geshkovski2025mathematical] for an extension to the (SA) dynamics.
@@
(i.e. all tokens lie in a common open hemisphere). Let $(x_{i}(t))_{i=1}^{n}$ be the solution of either the SA or USA dynamics with this initialization. Then there exist $x^{*}\in\mathbb{S}^{d-1}$ and positive constants $C,\lambda$ (depending on $n,\beta$ and the initialization) for which
@@
The analysis in [geshkovski2024dynamic] establishes the existence of exponentially long-lived metastable states. Starting from a well-separated initial configuration consisting of $k$ sets of tokens that are close relative to the distance to the set, the tokens first collapse within each set, forming $k$ tight groups during the time interval $0,T_{1}]$, where $T_{1}\sim\beta$. These clusters then persist on an interval $T_{1},T_{2}$ where $\log T_{2}\sim\beta$ before successive merging events occur. Geometrically, the flow of tokens remains near a manifold of $k$-cluster configurations, where the energy gradient is exponentially small and the motion of clusters is correspondingly slow.
@@
To build intuition, notice that as $\beta\to\infty$ the softmax velocity field in (SA) converges [^1] Formally, the limiting rule is $x_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\!\left(\mathop{\mathrm{arg\,max}}_{j}\langle x_{i}(t),x_{j}(t)\rangle\right)={\mathbf{P}}^{\perp}_{x_{i}(t)}x_{i}(t)=0\,.$ Following [hardmax], and to avoid this triviality, we forbid tokens from attending to themselves. to a hard $\mathop{\mathrm{arg\,max}}$ rule:
$\dot{x}_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\Bigl(\mathop{\mathrm{arg\,max}}_{j\neq i}\langle x_{i}(t),x_{j}(t)\rangle\Bigr)\,.$
Because $\beta$ is large, the (unique) closest pair $(\bar{\imath},\bar{\jmath})$—in the sense that
@@
—interacts on a timescale that is exponentially faster in $\beta$ than all other pairs. This leads to a deterministic merging event of that closest pair before any other interaction occurs.
@@
The dynamics (SA) are closely related to a continuous-time analogue of the classical Mean-Shift clustering algorithm. In its original form, the Mean-Shift algorithm defines clusters as basins of attraction of the modes of a kernel density estimator (KDE) [fukunaga1975estimation]. More precisely, given points $x_{1}(0),\ldots,x_{n}(0)\in\mathbb{R}^{d}$ drawn independently from a density $p$, and a kernel $K(\cdot)$ on $\mathbb{R}^{d}$ (typically Gaussian), recall that the KDE of $p$ is given by $\hat{p}=\hat{K}*\mu_{0}$, the convolution of $K$ with the empirical measure $\mu_{0}$ of the $x_{i}$; see, e.g., [tsybakov2008nonparametric, Chapter 1]. The following gradient-flow dynamics move each point toward the nearest mode of $K*\mu_{0}$:
@@
This analogy suggests tokens cluster into $K$ groups, where $K$ is the number of modes of the KDE $\hat{p}$. Using Edgeworth expansions and the Kac–Rice formula, we can show that $\mathbb{E}[K]$ is of order $\sqrt{\beta\log\beta}$ as $n\to\infty$, in the regime $n^{c}\leq\beta\leq n^{2-c}$ (with arbitrarily small fixed $c$), when $p$ is a Gaussian density on the real line [geshkovski2025numbermodesgaussiankernel]. Numerical simulations indicate that this order of magnitude should also hold on the circle, at least up to logarithmic factors. Recall, however, that this multiple-cluster state is metastable and will eventually collapse to a single one according to the results of Section 4.
@@
Consider an equiangular initialization [^2] Note that since $\rho_{0}\geq 0$ all points are initialized in the same hemisphere, as in Section 4.2. where $\langle x_{i}(0),x_{j}(0)\rangle=\rho_{0}\in[0,1]$, for all $i\neq j$. Because of symmetry, the configuration remains equiangular for all times under both SA and USA dynamics, and the entire system is characterized by the common correlation $\rho(t):=\langle x_{i}(t),x_{j}(t)\rangle$, for all $i\neq j$. The evolution of $\rho$ is described by a simple ordinary differential equation (ODE):
@@
$\dot{x}_{i}(t)=\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle\frac{x_{i}(t)}{\|x_{i}(t)\|},\frac{x_{j}(t)}{\|x_{j}(t)\|}\rangle}\,x_{j}(t),$
@@
while the magnitudes $r_{i}(t)$ satisfy a rule-dependent radial equation. Crucially, each normalization scheme induces a corresponding speed regulation factor $s_{i}(t)$ on token ii. For example, for Post-LN, taking $s_{i}(t)=1$ recovers (SA) while it can be shown that Pre-LN yields $s_{i}(t)=r_{i}(t)$ so that the directions of tokens with large magnitude are effectively slowed down. Other variants include Peri-LN [kim2025peri], Mix-LN [li2024mixln], nGPT [loshchilov2024ngpt], and sqrt-scaling [Nocisqrt]. All fit into this framework with different choices of $s_{i}(t)$.
@@
Under an equiangular initialization, we get $\langle\theta_{i}(t),\theta_{j}(t)\rangle=\rho(t)$ for all $i\neq j$ and $r_{i}(t)=r(t)$ for all tokens ii. The coupled ODEs governing $\rho(t)$ and $r(t)$ can be written explicitly for each scheme and solved explicitly as $t\to\infty$ using a simple linearization argument. It yields $1-\rho(t)\sim e^{-2t}$ for Post-LN while $1-\rho(t)\sim 1/t^{2}$ for Pre-LN. This marked difference in the rate of contraction to a single cluster—exponential vs. polynomial—confirms the practical wisdom that Pre-LN makes better use of depth by delaying contraction and hence avoiding representation collapse; see Figure 4.
@@
Theorem 7 ([chen2025clustering]).
@@
In particular, a single attention block already exhibits a phase transition: uniform contraction for subcritical $\gamma$, critical sparse mixing at $\gamma=\frac{1}{1-\rho}$, and an identity-like regime for supercritical $\gamma$. Since repeated layers amplify contraction multiplicatively, the one-step behavior fully determines long-time clustering. Thus the equiangular model provides a clean analytical description of how the logarithmic scaling $\beta_{n}\sim\log n$ stabilizes long-context attention by maintaining content-adaptive sparsity while avoiding collapse. Note also that a critical scaling of at $\beta\sim\log n$ also appears in perturbations of the exact equiangular model, in particular allowing dimension to be of order $\log n$ rather than $n$.
@@
$\mathrm{d}X_{i}(t)={\mathbf{P}}^{\perp}_{X_{i}(t)}\!\left(\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle X_{i}(t),X_{j}(t)\rangle}\,\mathrm{d}t\right)+\sqrt{2\kappa^{-1}}\,\mathrm{d}W_{i}(t),$
where $\kappa>0$ controls the relative strength of the stochastic and drift terms, and $W_{1},\dots,W_{n}$ are independent Brownian motions on $\mathcal{S}^{d-1}$. In the limit $\kappa\to\infty$, one recovers the deterministic dynamics (USA).
@@
$\mathrm{d}X(t)={\mathbf{P}}^{\perp}_{X(t)}\!\left(\int e^{\beta\langle X(t),y\rangle}\,\mu_{t}(\mathrm{d}y)\right)\mathrm{d}t+\sqrt{2\kappa^{-1}}\,\mathrm{d}W(t),$ (8)
@@
$\partial_{t}\mu_{t}+\kappa^{-1}\Delta\mu_{t}=\int e^{\beta\langle x,y\rangle}y\,\mu_{t}(\mathrm{d}y),$ (9)

### v2

@@ Abstract.
We develop a mathematical framework that interprets Transformer attention as an interacting particle system and studies its continuum (mean-field) limits. By idealizing attention on the sphere, we connect Transformer dynamics to Wasserstein gradient flows, synchronization models (Kuramoto), and mean-shift clustering. Central to our results is a global clustering phenomenon whereby tokens cluster asymptotically after long metastable states where they are arranged into multiple clusters. We further analyze a tractable equiangular reduction to obtain exact clustering rates, show how commonly used normalization schemes alter contraction speeds, and identify a phase transition for long-context attention. The results highlight both the mechanisms that drive representation collapse and the regimes that preserve expressive, multi-cluster structure in deep attention architectures.
@@
A standard Transformer is a stack of attention blocks, each consisting of two main components: an attention layer and a feed-forward MLP layer. Both components are followed by normalization and residual connections. Schematically, if $\mathbf{X}=(X_{1},\ldots,X_{n})\in\mathbb{R}^{d\times n}$ denotes the matrix of $n$ tokens of dimension $d$, a single layer performs the composition of the maps:
$X_{i}$ $\longmapsto X_{i}+\mathrm{Attention}(\mathbf{X})_{i},$
$X_{i}$ $\longmapsto X+\mathrm{MLP}(X_{i})\,$,
for $i=1,\ldots,n$. In particular, all the tokens interact in the Attention layer but not in the MLP layer. Normalization steps are inserted before or after each layer depending on the implementation; see Section 6.2 for some examples. These residual connections, together with layer normalization [ba2016layer, zhang2019root], play a crucial stabilizing role, analogous to skip connections in residual networks [he2016identity, haber2017stable]. We will later see that such normalization can be idealized as a projection onto the unit sphere, a simplification that preserves essential dynamical features of trained models.
@@
$\mathrm{Attention}(\mathbf{X})_{i}=\sum_{j=1}^{n}\frac{\exp\!\big(\beta\langle QX_{i},KX_{j}\rangle\big)}{\sum_{k=1}^{n}\exp\!\big(\beta\langle QX_{i},KX_{k}\rangle\big)}\,VX_{j}.$ (1)
Equation (1) can thus be interpreted as a nonlinear interaction rule in which each token updates as a weighted average of all others, with weights depending on their similarity in feature space. This operator will later serve as the starting point for the simplified continuous-time models studied in this paper.
@@
Modern variants of the Transformer introduce additional components: hierarchical structures for vision tasks [liu2021swin], sparse or long-range attention [beltagy2020longformer, child2019generating], low-rank adaptation mechanisms [hu2022lora], and mixture-of-experts [shazeer2017outrageously, boixadsera2025powerfinegrainedexpertsgranularity] architectures that route tokens through specialized subnetworks. These extensions have led to dramatic empirical gains but lie outside the scope of this paper. Our goal is to isolate and analyze the mathematical core of the architecture, the attention mechanism itself, and to understand how it regulates token dynamics.
@@
$\mathcal{E}_{\beta}(\mu)=\frac{1}{2\beta}\iint e^{\beta\langle x,y\rangle}\,\mathrm{d}\mu(x)\mathrm{d}\mu(y).$ (3)
For (SA), it also corresponds to a gradient flow over the space of probability measures, albeit with respect to a different (Hessian) metric [geshkovski2025mathematical, Li21]. In the large-$\beta$ regime, these two dynamics recover familiar PDEs at leading order. Indeed the continuity equation associated with (SA) converges formally to a reverse heat equation. This anti-diffusive limit foreshadows the clustering behavior described in the next section. In contrast, for (USA) with appropriate time rescaling, it converges to a porous medium equation; see [bruno2025multiscale].
The (SA) and (USA) dynamics are minimal but faithful abstractions of self-attention. They retain its essential nonlinear and nonlocal features while remaining amenable to analysis. As discussed in the next section, both exhibit clustering effects similar to those observed in trained Transformers, providing a tractable framework to study attention-driven representation dynamics. In fact, when restricted to the circle, these dynamics reduce to variants of the Kuramoto model, the classical framework to study synchronization which corresponds precisely to the clustering behavior alluded to above.
3.3. The Kuramoto connection.
@@
We now turn to one of the most striking properties of the dynamics (USA) and (SA): the spontaneous emergence of clusters. In fact, clustering can be observed in trained Transformer models: Figure 1 indicates token embeddings exhibit a progressive concentration of pairwise inner products near $1$, revealing the gradual formation of clusters.
Figure 1. Histograms of pairwise inner products $\{\langle x_{i}(t),x_{j}(t)\rangle\}_{i\neq j}$ at layers $t=0$, 55, 3030, and $48$ in the pre-trained ALBERT XLarge v2 model [lanalbert] available on Huggingface. The progressive concentration of mass near 11 across layers illustrates the emergence of clustering in token embeddings. Figure reproduced from [geshkovski2025mathematical].
@@
The result follows from the general theory of consensus dynamics on compact manifolds [markdahl2017almost]. Since the dynamics (SA) are a smooth gradient flow of an analytic energy functional, the classical theorem of Łojasiewicz [lojasiewicz1963propriete] guarantees that every trajectory converges to a stationary point. A detailed analysis of the critical set shows that, except for the fully synchronized configuration, all other stationary points are saddles by exhibiting an escape direction. This result also holds for weighted tokens that evolve according to the Wasserstein-Fisher-Rao gradient flow whereby particles evolve and are also dynamically reweighted [chen2025clustering]. By the center-stable manifold theorem [shub2013global, Thm. III.7], the set of initial conditions whose trajectories converge to such saddles is contained in a countable union of lower-dimensional manifolds and thus has measure zero. Consequently, almost every initialization leads to convergence toward the synchronized equilibrium. This argument, first established in [markdahl2017almost] and subsequently refined in [criscitiello2024synchronization], provides a complete characterization of the long-time behavior of the (USA) dynamics; see [geshkovski2025mathematical] for an extension to the (SA) dynamics.
@@
(i.e. all tokens lie in a common open hemisphere). Let $(x_{i}(t))_{i=1}^{n}$ be the solution of either the (SA) or (USA) dynamics with this initialization. Then there exist $x^{*}\in\mathbb{S}^{d-1}$ and positive constants $C,\lambda$ (depending on $n,\beta$ and the initialization) for which
@@
The analysis in [geshkovski2024dynamic] establishes the existence of exponentially long-lived metastable states. Starting from a well-separated initial configuration consisting of $k$ sets of tokens that are close together relative to other sets, the tokens first collapse within each set, forming $k$ tight groups during the time interval $0,T_{1}]$, where $T_{1}\sim\beta$. These clusters then persist on an interval $T_{1},T_{2}]$ where $\log T_{2}\sim\beta$ before successive merging events occur. Geometrically, the flow of tokens remains near a manifold of $k$-cluster configurations, where the energy gradient is exponentially small and the motion of clusters is correspondingly slow.
@@
To build intuition, notice that as $\beta\to\infty$ the softmax velocity field in (SA) converges [^1] Formally, the limiting rule is $x_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\!\left(\mathop{\mathrm{arg\,max}}_{x_{j}(t)}\langle x_{i}(t),x_{j}(t)\rangle\right)={\mathbf{P}}^{\perp}_{x_{i}(t)}x_{i}(t)=0\,.$ Following [hardmax], and to avoid this triviality, we forbid tokens from attending to themselves. to a hard $\mathop{\mathrm{arg\,max}}$ rule:
$\dot{x}_{i}(t)={\mathbf{P}}^{\perp}_{x_{i}(t)}\Bigl(\mathop{\mathrm{arg\,max}}_{x_{j}(t)\neq x_{i}(t)}\langle x_{i}(t),x_{j}(t)\rangle\Bigr)\,.$
Because $\beta$ is large, the (unique) closest pair $(\bar{\imath},\bar{\jmath})$ in the sense that
@@
interacts on a timescale that is exponentially faster in $\beta$ than all other pairs. This leads to a deterministic merging event of that closest pair before any other interaction occurs.
@@
The dynamics (SA) are closely related to a continuous-time analogue of the classical Mean-Shift clustering algorithm. In its original form, the Mean-Shift algorithm defines clusters as basins of attraction of the modes of a kernel density estimator (KDE) [fukunaga1975estimation]. More precisely, given points $x_{1}(0),\ldots,x_{n}(0)\in\mathbb{R}^{d}$ drawn independently from a density $p$, and a kernel $K(\cdot)$ on $\mathbb{R}^{d}$ (typically Gaussian), recall that the KDE of $p$ is given by $\hat{p}=K*\mu_{0}$, the convolution of $K$ with the empirical measure $\mu_{0}$ of the $x_{i}$; see, e.g., [tsybakov2008nonparametric, Chapter 1]. The following gradient-flow dynamics move each point toward the nearest mode of $K*\mu_{0}$:
@@
This analogy suggests tokens cluster into $M$ groups, where $M$ is the number of modes of the KDE $\hat{p}$. Using Edgeworth expansions and the Kac–Rice formula, we can show that $\mathbb{E}[M]$ is of order $\sqrt{\beta\log\beta}$ as $n\to\infty$, in the regime $n^{c}\leq\beta\leq n^{2-c}$ (with arbitrarily small fixed $c$), when $p$ is a Gaussian density on the real line [geshkovski2025numbermodesgaussiankernel]. Numerical simulations indicate that this order of magnitude should also hold on the circle, at least up to logarithmic factors. Recall, however, that this multiple-cluster state is metastable and will eventually collapse to a single one according to the results of Section 4.
@@
Consider an equiangular initialization [^2] Note that since $\rho_{0}\geq 0$ all points are initialized in the same hemisphere, as in Section 4.2. where
$\langle x_{i}(0),x_{j}(0)\rangle=\rho_{0}\in[0,1]\quad\forall\ i\neq j.$
Because of symmetry, the configuration remains equiangular for all times under both (SA) and (USA) dynamics, and the entire system is characterized by the common correlation $\rho(t):=\langle x_{i}(t),x_{j}(t)\rangle$, for all $i\neq j$. The evolution of $\rho$ is described by a simple ordinary differential equation (ODE):
@@
$\dot{x}_{i}(t)=\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle\frac{x_{i}(t)}{\|x_{i}(t)\|},\frac{x_{j}(t)}{\|x_{j}(t)\|}\rangle}\,\frac{x_{j}(t)}{\|x_{j}(t)\|},$
@@
while the magnitudes $r_{i}(t)$ satisfy a rule-dependent radial equation. Crucially, each normalization scheme induces a corresponding speed regulation factor $s_{i}(t)$ on token $i$. For example, for Post-LN, taking $s_{i}(t)=1$ recovers (SA) while it can be shown that Pre-LN yields $s_{i}(t)=r_{i}(t)$ so that the directions of tokens with large magnitude are effectively slowed down. Other variants include Peri-LN [kim2025peri], Mix-LN [li2024mixln], nGPT [loshchilov2024ngpt], and sqrt-scaling [Nocisqrt]. All fit into this framework with different choices of $s_{i}(t)$.
@@
Under an equiangular initialization, we get $\langle\theta_{i}(t),\theta_{j}(t)\rangle=\rho(t)$ for all $i\neq j$ and $r_{i}(t)=r(t)$ for all tokens $i$. The coupled ODEs governing $\rho(t)$ and $r(t)$ can be written explicitly for each scheme and solved explicitly as $t\to\infty$ using a simple linearization argument. It yields $1-\rho(t)\sim e^{-2t}$ for Post-LN while $1-\rho(t)\sim 1/t^{2}$ for Pre-LN. This marked difference in the rate of contraction to a single cluster—exponential vs. polynomial—confirms the practical wisdom that Pre-LN makes better use of depth by delaying contraction and hence avoiding representation collapse; see Figure 4.
@@
Theorem 7 ([chen2025quantitative]).
@@
In particular, a single attention block already exhibits a phase transition: uniform contraction for subcritical $\gamma$, critical sparse mixing at $\gamma=\frac{1}{1-\rho}$, and an identity-like regime for supercritical $\gamma$. Since repeated layers amplify contraction multiplicatively, the one-step behavior fully determines long-time clustering. Thus the equiangular model provides a clean analytical description of how the logarithmic scaling $\beta_{n}\sim\log n$ stabilizes long-context attention by maintaining content-adaptive sparsity while avoiding collapse. Note also that a critical scaling of order $\beta\sim\log n$ also appears in perturbations of the exact equiangular model, in particular allowing dimension to be of order $\log n$ rather than $n$.
@@
$\mathrm{d}X_{i}(t)={\mathbf{P}}^{\perp}_{X_{i}(t)}\!\left(\frac{1}{n}\sum_{j=1}^{n}e^{\beta\langle X_{i}(t),X_{j}(t)\rangle}X_{j}(t)\,\mathrm{d}t\right)+\sqrt{2\kappa^{-1}}\,\mathrm{d}W_{i}(t),$
where $\kappa>0$ controls the relative strength of the stochastic and drift terms, and $W_{1},\dots,W_{n}$ are independent Brownian motions on $\mathbb{S}^{d-1}$. In the limit $\kappa\to\infty$, one recovers the deterministic dynamics (USA).
@@
$\mathrm{d}X(t)={\mathbf{P}}^{\perp}_{X(t)}\!\left(\int e^{\beta\langle X(t),y\rangle}\,y\mu_{t}(\mathrm{d}y)\right)\mathrm{d}t+\sqrt{2\kappa^{-1}}\,\mathrm{d}W(t),$ (8)
@@
$\partial_{t}\mu_{t}+\kappa^{-1}\Delta\mu_{t}=\nabla\cdot\left(\mu_{t}\int e^{\beta\langle\cdot,y\rangle}y\,\mu_{t}(\mathrm{d}y)\right),$ (9)